The Chunnel connects the United Kingdom and France by a railway tunnel under the English Channel. The British started tunneling 2.5 months before the French and averaged 0.63 kilometers per month. The French averaged 0.47 kilometers per month. When the two sides met, they had tunneled 37.9 kilometers. How many kilometers of tunnel did each country build? If the French started tunnelingon February 28, 1988, approximately when did the two sides meet?: The Chunnel connects the United Kingdom and France by a railway tunnel under the English Channel. The British started tunneling 2.5 months before the French and averaged 0.63 kilometers per month. The French averaged 0.47 kilometers per month. When the two sides met, they had tunneled 37.9 kilometers. How many kilometers of tunnel did each country build? If the French started tunnelingon February 28, 1988, approximately when did the two sides meet?
Answer by ankor@dixie-net.com(4531)

The Channel connects the United Kingdom and France by a railway tunnel under the English Channel. The British started tunneling 2.5 months before the French and averaged 0.63 kilometers per month. The French averaged 0.47 kilometers per month. When the two sides met, they had tunneled 37.9 kilometers. How many kilometers of tunnel did each country build? If the French started tunneling on February 28, 1988, approximately when did the two sides meet?
:
Let t = months tunneling by the British
then
(t-2.5) = months tunneling by the French
;
Write a distance equation: Dist = speed * time
:
British dist + French dist = 37.9
.63t + .47(t-2.5) = 37.9
:
.63t + .47t - 1.175 = 37.9
:
.63t + .47t = 37.9 + 1.175
:
1.1t = 39.075
t = 39.075/1.1
t = 35.52 months
then
35.52 * .63 = 22.38 km Built by the British
And
(35.52-2.5) * .47 = 15.52 km built by the French
:
French construction time 35.52 - 2.5 = 33.02 months
:
Using an average of 30.4 days per month: 30.4 * 33.02 ~ 1004 days
:
1004/365 = 2.75 yrs, or 2 yrs & 9 months
:
Starting it Feb 28, 1988, then it would be about Dec 1, 1990